The head of the PE Department of a certain high school claims that the mean height ofGrade 7 students is 163 cm. The mean height of 45 randomly selected Grade 7 studentsis 161 cm. Using 0.01 significance level, can it be concluded that the mean height ofGrade 7 students is different from 163 cm as claimed by the Head of the PE Department?

1
Expert's answer
2022-05-27T00:18:25-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=163"

"H_a:\mu\not=163"

This corresponds to a two-tailed test, for which a t-test for two population means, with two independent samples, with unknown population standard deviations will be used.

Based on the information provided, the significance level is "\alpha = 0.01," and the degrees of freedom are "df = 45-1=44"

Hence, it is found that , using t-table, the critical value for this two-tailed test is "\:\:t_c\:=2.69" for "\alpha =0.01" and "\:\:df=44" .

The rejection region for this two-tailed test is "R=\left\{t:∣t∣>2.69\right\}."

The z-statistic is computed as follows:



"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{161-163}{\sqrt{45}}\approx-0.3"

Since it is observed that "z = -0.066> -2.3263=z_c," it is then concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is less than 89, at the "\alpha = 0.01" significance level.

Since it is observed that "\u2223t\u2223=0.3<t \nc\n\u200b\t\n =2.69" it is then concluded that the null hypothesis is not rejected.